Calculate the number of photon emitted per second by a10 W sodium vapour lamp. Assume that 60% of the consumed energy if converted into light. Wavelength of sodium light = 590 nm.

Grade:11

1 Answers

Jitender Pal
9 years ago
Sol. P = 10 W ∴ E in 1 sec = 10 J % used to convert into photon = 60% ∴ Energy used = 6 J Energy used to take out 1 photon = hc/λ = 6.63* 10^-34 * 3 * 10^8/590 * 10^-9 = 6.633/590 * 10^-17 No. of photons used = 6/6.63 *3/590 * 10^-17 = 6* 590/6.63 * 3*10^17 = 176.9 * 10^17 = 1.77 * 10^19